An angle measure 20 degrees more than 4 times the measure of its complementary angle. Find the measures of both angles.
Answer by waynest(281)

angle 1 = x
angle 2 = 4x + 20
x + 4x + 20 = 90
5x + 20 - 20 = 90 - 20
5x = 70
5x/5 = 70/5
x = 14
check:
14 + (4(14) + 20) = 90
14 + (56 + 20) = 90
14 + 76 = 90
90 = 90